# MLLM-CL: Continual Learning for Multimodal Large Language Models

### v1

@@ 4.2 Router Selection with MLLM
Framework of MR-LoRA. During inference, with expert learning and router selection, the overall framework of the proposed method is illustrated in Fig. 5. Our MR-LoRA performs two-stage inference for a given multimodal input, consisting of a routing phase followed by a prediction phase. In the first stage, the expert selection router is performed to select a domain or ability-specific expert. Then, the selected expert is combined with the pre-trained backbone to output the final response. On the one hand, by decoupling the learning of different domains or abilities, we avoid potential distribution conflict and can learn a good expert for a given task. On the other hand, the proposed router selection strategy largely explores the advantages of MLLMs to improve the flexibility and accuracy of expert selection, ensuring promising final prediction performance during continual learning.
@@ 5.1 Experimental Setup
Evaluation Metric. We report Last and Average following standard metrics in continual learning l2p ; guo2025hide . Last is the accuracy of all seen tasks after learning the last task. Average is the average accuracy of each task during the training process, i.e., $\textit{Average}=\sum_{i=1}^{t}\text{acc}_{i},$ where $t$ is the task that the model is learning, $\text{acc}_{i}$ is the accuracy of the $i$-th previous learned task.
@@
Table 2: Results for domain continual learning in MLLM-CL benchmark. denotes the original method with replay data.
Method RS (%) Med (%) AD (%) Sci (%) Fin (%)
Last Average Last Average Last Average Last Average Last Average
LLaVA-v1.5-7b liu2023llava 32.29 - 28.28 - 15.59 - 35.55 - 62.56 -
LoRA hu2021lora 26.75 80.72 25.76 59.68 0.79 40.51 18.69 18.64 70.44 28.49
LoRA hu2021lora 29.57 80.87 29.19 58.60 7.09 38.95 19.55 36.41 63.60 36.78
MoELoRA chen2024coin 21.42 80.05 25.29 57.26 0.79 37.03 17.01 19.65 60.34 24.97
MoELoRA chen2024coin 40.23 80.00 23.58 56.91 5.19 34.69 18.35 31.70 74.89 31.36
O-LoRA wang2023orthogonal 62.68 80.22 35.17 67.56 16.93 51.51 34.44 44.28 92.16 48.28
O-LoRA wang2023orthogonal 76.21 80.13 51.34 70.23 36.50 61.35 42.64 53.34 90.20 59.38
L2P l2p 63.82 80.02 34.63 68.86 22.96 51.57 38.58 45.12 92.98 50.59
L2P l2p 75.21 80.09 38.50 68.64 32.31 54.79 41.05 48.68 88.05 55.02
ModalPrompt zeng2024modalprompt 65.99 80.11 37.35 59.66 23.27 46.86 37.61 42.97 87.60 50.36
ModalPrompt zeng2024modalprompt 64.77 80.11 38.60 60.99 20.61 50.67 29.98 41.97 88.22 48.44
HiDe-LLaVA guo2025hide 41.17 80.91 30.33 65.47 18.73 39.78 37.08 32.92 92.21 43.90
HiDe-LLaVA guo2025hide 75.36 81.51 39.23 62.37 37.17 49.37 45.02 50.61 81.89 55.73
Ours 79.87 80.82 62.71 72.19 51.89 65.41 52.48 62.52 89.69 67.31
@@
Table 3: Results for ability continual learning in MLLM-CL benchmark. denotes the original method with replay data.
Method OCR (%) M & L (%) VP (%) GUI Agent (%)
Last Average Last Average Last Average Last Average
LLaVA-v1.5-7b liu2023llava 31.10 - 25.60 - 60.91 - 25.00 -
LoRA hu2021lora 10.80 31.10 26.40 15.25 27.16 39.38 30.83 23.80
LoRA hu2021lora 12.00 31.10 22.50 17.20 14.34 40.03 33.33 20.54
MoELoRA chen2024coin 15.10 30.60 28.30 16.40 11.80 38.76 27.78 20.75
MoELoRA chen2024coin 11.90 30.60 24.70 15.80 23.10 38.78 23.75 20.86
O-LoRA wang2023orthogonal 18.50 32.90 26.10 32.85 51.90 44.16 32.50 32.25
O-LoRA wang2023orthogonal 28.60 32.90 32.80 32.95 51.40 43.67 32.50 36.33
L2P l2p 25.50 31.40 32.00 31.75 48.22 41.59 16.25 30.39
L2P l2p 25.50 32.30 26.10 29.75 61.80 40.11 18.47 32.97
ModalPrompt zeng2024modalprompt 31.00 31.00 31.80 28.20 61.68 41.83 23.75 37.06
ModalPrompt zeng2024modalprompt 31.00 31.00 24.40 27.70 60.91 38.77 21.25 34.39
HiDe-LLaVA guo2025hide 21.70 31.60 30.10 32.45 63.96 37.40 25.00 35.19
HiDe-LLaVA guo2025hide 24.20 30.40 27.20 33.90 44.16 40.52 34.17 32.43
Ours 35.00 35.00 33.10 33.10 62.06 43.35 29.31 39.87
@@ 5.2 Results and Analysis
Domain Continual Learning. Results on domain continual learning are shown in Tab. 2. We can find that our MR-LoRA significantly outperforms existing continual learning methods in both Last and Average accuracy and achieves over 15% gains compared with most baselines. Notably, our performance in the science domain is higher than training from scratch individually (52.48 vs. 50.17), which means our MLLM based router selects a more suitable expert instead of the science expert to solve some questions and this expert can handle these questions better than the original science expert. Fig. 6 shows an example from the science task where the science expert fails to solve the problem while the medical expert succeeds. Moreover, we observe that task decoupling (i.e. parameter isolation) methods like L2P l2p , HiDe-LLaVA guo2025hide , O-LoRA wang2023orthogonal achieve better performance compared with task non-decoupling methods like MoELoRA chen2024coin .This means the domain gap between different tasks is big and introducing new learnable modules in each task can alleviate more catastrophic forgetting. Even in the case of IID testing, the larger domain gap improves the challenge of the benchmark, which shows the effectiveness and necessity of our proposed benchmark.
Figure 6: An example demonstrating that the medical expert handles certain science questions better than the original science expert. Introducing our MLLM-enhanced router enables automatic selection of the most appropriate experts.
Ability Continual Learning. Tab. 3 shows the results of ability continual learning. In this setting, MR-LoRA still outperforms all baselines and improves the performance in all abilities. However, most baselines suffer from severe catastrophic forgetting and the performance in previous tasks is worse than that in the original model (i.e., LLaVA-v1.5-7b), which shows the limitation of existing continual learning methods in real-world applications. The same as domain continual learning, we can observe that the performance of math & logic in MR-LoRA is higher than training an individual LoRA from scratch (33.1 vs. 33.0), which reflects the flexibility and rationality of MR-LoRA in router selection. This can be viewed as knowledge transfer within our MLLM-enhanced router.
Rank of Expert LoRA. From the results in Tab. 4, we find that our method performs well even at very low ranks (e.g., 8), demonstrating its parameter efficiency. This indicates that even if the number of tasks to be learned is large, our method can still achieve a good performance with only a small increase in parameters. Besides, as the expert rank increases, performance can continue to improve slightly because of more trainable parameters. We choose 128 as the default rank following the default LoRA finetuning setting for LLaVA liu2023llava .
Table 4: Ablation study of LoRA rank for each expert LoRA in domain continual learning. We show the last accuracy below.
@@
8 79.03 60.05 51.58 47.51 88.24
16 79.19 60.46 52.13 48.31 88.75
32 79.32 61.20 51.95 49.33 88.84
64 79.67 62.16 51.87 50.98 89.25
128 79.87 62.74 51.89 52.48 88.69
256 79.92 63.32 52.20 54.53 90.17
Router Accuracy. We ablate the number of samples for routing data and report the router selection accuracy and last accuracy in domain and ability continual learning. The results are shown in Secs. 5.2 and 6. In domain continual learning, we find that our method can achieve an excellent performance (over 96% in average) using only 20 samples to train the router, which means our method closes the gap of training each task individually. Note that the number of samples we used is much smaller than the number of training samples (60k). Besides, with more sampling data, the router selection accuracy improves and the performance of MR-LoRA slightly increases. In ability continual learning, the performance of  MR-LoRA  achieves satisfactory performance when the shot of router tuning is 10. It is interesting that the router accuracy of the OCR task is around 40%, but our method can achieve a comparable even better performance (at least over 34) compared with directly finetuning an OCR LoRA expert (35.0). This means MR-LoRA use other experts to solve the OCR task and these experts perform well on these test samples. It is reasonable that OCR is a basic and fundamental ability that the math, GUI Agent experts are also able to extract equations and web texts in the images.
Visualization. Fig. 7 provides examples during DCL and ACL, respectively. We can find that some baselines like LoRA hu2021lora , MoELoRA chen2024coin , HiDe-LLaVA guo2025hide overfit to the last learned task and output options that do not exist in domain continual learning. There are also some baselines that do not understand the problem correctly, such as L2P l2p , ModalPrompt zeng2024modalprompt in domain continual learning and HiDe-LLaVA guo2025hide , O-LoRA wang2023orthogonal , MoELoRA chen2024coin in ability continual learning.
Figure 7: Visualization of MR-LoRA and other baselines under domain continual learning and ability continual learning. The left part is testing autonomous driving task after learning all domain tasks, while the right part is testing OCR tasks after learning all ability tasks.
@@
# Replay Samples Router Accuracy (%) Last Accuracy (%)
RS (%) Med (%) AD (%) Sci (%) Fin (%) RS (%) Med (%) AD (%) Sci (%) Fin (%)
100 99.71 93.06 100 98.70 100 80.69 62.24 52.49 54.37 89.69
50 99.47 98.60 100 97.30 100 80.64 63.28 52.49 53.99 89.69
30 99.95 97.66 100 94.98 99.98 80.80 63.18 52.49 53.71 89.68
20 97.83 95.85 96.82 91.86 100 79.87 62.71 51.89 52.48 89.69
10 98.91 69.20 98.54 93.90 100 80.25 54.32 52.21 53.11 89.69
Table 6: Router accuracy under different amount of replay data in ability continual learning. The left part is the router selection accuracy and the right part is task accuracy after learning the last task.
# Replay Samples Router Accuracy (%) Last Accuracy (%)
OCR (%) M&L (%) VP (%) GUI Agent (%) OCR (%) M&L (%) VP (%) GUI Agent (%)
100 39.70 97.50 100 100 34.10 32.70 62.06 29.31
50 45.40 97.80 99.87 100 34.70 32.70 61.93 29.31
30 49.10 99.10 99.11 100 34.80 33.10 61.42 29.31
20 50.00 97.80 100 100 35.00 33.10 62.06 29.31
10 33.90 94.80 100 100 35.00 32.70 62.06 29.31
@@ 6 Conclusion
In this paper, we first propose MLLM-CL benchmark, a novel benchmark including domain continual learning and ability continual learning. In domain continual learning, we select five specific domains (remote sensing, medical, science, autonomous driving, and finance) and focus on IID evaluation. In ability continual learning, we consider a more practical setting where the training and test sets are non-IID. We select four common and fundamental abilities for MLLM to learning sequentially: OCR, math & logic, visual perception and GUI agent. To solve the two settings in the MLLM-CL benchmark, we first analyze the task conflict between different tasks and then propose an MLLM enhanced router selection method MR-LoRA . Comprehensive experiments and analyses validate the necessity of our MLLM-CL benchmark and show the effectiveness and efficiency of our proposed method. We believe that our carefully designed benchmark and MR-LoRA  can serve as a foundation for continual learning in multimodal large language models and will introduce an innovative and practical direction of continual learning and MLLM to the community.

### v2

@@ 4.2 Inference: Router Selection with MLLM
Framework of MR-LoRA. During inference, with expert learning and router selection, the overall framework of the proposed method is illustrated in Fig.˜5. Our MR-LoRA performs two-stage inference for a given multimodal input, consisting of a routing phase followed by a prediction phase. In the first stage, the expert selection router is performed to select a domain or ability-specific expert. Then, the selected expert is combined with the pre-trained backbone to output the final response. On the one hand, by decoupling the learning of different domains or abilities, we avoid potential distribution conflict and can learn a good expert for a given task. On the other hand, the proposed router selection strategy largely explores the advantages of MLLMs to improve the flexibility and accuracy of expert selection, ensuring promising final prediction performance during continual learning. The proposed MLLM-based routing mechanism offers notable advantages: (1) The MLLM’s strong multimodal understanding capacity ensures robust expert selection performance on complex multimodal inputs. (2) The selection router is parameter-efficient and learned with few-shot unlabeled image-question pairs, allowing on-the-fly adaptation.
@@ 5.1 Experimental Setup
Evaluation Metric. We report the last accuracy, which is the accuracy of all seen tasks after learning the last task, mean finetune accuracy (MFT), mean final accuracy (MFN), mean average accuracy (MAA), and backward transfer (BWT) following standard metrics in continual learning (Guo et al., 2025a; Chen et al., 2025). The detailed calculation of each metric is shown in the Sec.˜A.3.
@@
Table 2: Results for LLaVA-based domain continual learning in MLLM-CL benchmark. denotes the original method with replay data.
[…]
Zeroshot 32.29 28.28 15.59 35.55 62.56 34.85 - - -
Oracle 81.06 65.83 54.17 56.86 91.14 69.81 - - -
LoRA-FT (Hu et al., 2021) 69.65 41.59 25.43 40.88 87.45 64.98 53.00 61.13 -14.97
LoRA-FT (Hu et al., 2021) 76.54 50.27 43.01 43.32 89.85 66.32 60.60 64.72 -7.15
O-LoRA (Wang et al., 2023a) 74.64 44.42 30.02 41.47 87.15 65.16 55.54 62.12 -12.03
O-LoRA (Wang et al., 2023a) 76.94 41.17 34.18 39.61 83.22 60.49 55.02 60.73 -6.83
MoELoRA (Chen et al., 2024a) 77.54 41.85 27.62 40.13 86.75 64.94 54.78 61.76 -12.70
MoELoRA (Chen et al., 2024a) 77.63 49.54 39.08 41.04 89.21 66.24 59.30 64.81 -8.68
CL-MoE (Huai et al., 2025) 71.34 46.84 26.33 41.17 88.74 66.06 54.88 61.79 -13.96
CL-MoE (Huai et al., 2025) 76.58 52.31 39.65 45.64 90.21 66.65 60.88 64.95 -7.22
HiDe (Guo et al., 2025a) 74.31 48.95 33.21 38.54 81.55 60.77 55.31 60.68 -6.82
HiDe (Guo et al., 2025a) 74.80 42.29 34.03 38.01 79.22 60.83 53.67 61.81 -8.95
SEFE (Chen et al., 2025) 77.26 50.37 37.21 40.87 86.82 65.01 58.51 63.63 -8.13
SEFE (Chen et al., 2025) 78.43 52.85 46.21 47.76 89.33 66.89 62.92 66.51 -4.97
DISCO (Guo et al., 2025b) 76.03 45.20 43.79 42.33 88.95 64.43 59.26 63.35 -6.46
DISCO (Guo et al., 2025b) 77.78 46.25 50.45 49.51 89.71 65.27 62.74 64.92 -3.17
MR-LoRA (Ours) 80.87 65.32 54.12 56.71 91.12 69.64 69.63 71.06 -0.01
@@
Table 3: Results for LLaVA-based ability continual learning in MLLM-CL benchmark.
[…]
Zeroshot 31.20 30.20 60.79 10.00 33.05 - - -
Oracle 33.60 36.50 65.10 32.50 41.93 - - -
LoRA-FT (Hu et al., 2021) 23.60 33.70 55.84 32.50 41.28 36.41 36.58 -6.49
LoRA-FT (Hu et al., 2021) 21.80 32.70 58.38 28.75 40.32 35.41 36.32 -6.55
O-LoRA (Wang et al., 2023a) 29.60 32.90 52.41 33.75 39.72 37.16 35.42 -3.41
O-LoRA (Wang et al., 2023a) 29.60 31.30 60.79 27.50 39.96 37.30 36.34 -3.55
MoELoRA (Chen et al., 2024a) 26.70 32.80 56.85 27.22 39.45 35.89 36.07 -4.75
MoELoRA (Chen et al., 2024a) 19.80 32.20 54.19 30.00 40.35 34.05 35.39 -8.41
CL-MoE (Huai et al., 2025) 19.90 32.70 53.43 30.69 40.50 34.18 35.65 -8.43
CL-MoE (Huai et al., 2025) 25.40 31.80 60.91 30.00 41.22 37.03 37.28 -5.59
HiDe (Guo et al., 2025a) 24.60 32.10 46.32 28.75 37.98 32.94 34.60 -6.72
HiDe (Guo et al., 2025a) 24.60 28.40 30.71 23.75 36.84 26.86 33.54 -13.30
SEFE (Chen et al., 2025) 26.00 33.40 57.74 33.75 40.98 37.72 36.59 -4.35
SEFE (Chen et al., 2025) 25.60 34.80 57.61 31.39 42.25 37.35 37.93 -6.53
DISCO (Guo et al., 2025b) 32.90 33.10 60.15 30.14 39.02 39.07 36.57 0.07
DISCO (Guo et al., 2025b) 34.20 35.00 61.55 27.50 40.14 39.56 37.85 -0.77
MR-LoRA (Ours) 33.70 36.20 65.10 32.50 41.89 41.88 38.86 -0.02
Table 4: Results for InternVL-based domain continual learning in MLLM-CL benchmark. denotes the original method with replay data.
[…]
Zeroshot 31.16 29.81 14.06 33.93 64.32 34.66 - - -
Oracle 81.49 66.42 54.56 54.48 91.24 69.64 - - -
LoRA-FT (Hu et al., 2021) 69.93 52.17 33.04 42.67 91.07 69.06 57.78 65.22 -14.11
LoRA-FT (Hu et al., 2021) 77.06 47.55 42.67 43.31 91.44 69.43 60.41 67.45 -11.28
MoELoRA (Chen et al., 2024a) 69.90 52.08 33.17 42.19 90.58 68.83 57.58 65.97 -14.06
MoELoRA (Chen et al., 2024a) 76.74 52.65 38.81 42.15 89.84 67.90 60.04 66.01 -9.83
HiDe (Guo et al., 2025a) 75.40 57.66 36.73 41.48 88.59 65.26 59.97 65.94 -6.60
HiDe (Guo et al., 2025a) 53.17 52.61 40.85 47.04 89.17 64.20 56.57 61.06 -9.54
DISCO (Guo et al., 2025b) 75.12 50.69 52.41 50.67 90.86 68.85 63.95 68.14 -6.12
DISCO (Guo et al., 2025b) 77.90 47.50 49.13 49.37 90.92 68.55 62.96 67.81 -6.98
MR-LoRA (Ours) 81.48 65.80 54.56 54.40 91.07 69.51 69.46 71.27 -0.06
Table 5: Results for InternVL-based ability continual learning in MLLM-CL benchmark.
[…]
Zeroshot 30.00 31.20 56.09 2.50 29.95 - - -
Oracle 32.20 33.40 67.77 33.75 41.78 - - -
LoRA-FT (Hu et al., 2021) 21.40 32.80 60.28 29.86 40.84 36.08 36.38 -6.35
LoRA-FT (Hu et al., 2021) 26.30 34.20 62.56 31.25 41.63 38.58 37.38 -4.07
O-LoRA (Wang et al., 2023a) 25.50 32.30 64.59 24.44 38.64 36.71 36.05 -2.57
O-LoRA (Wang et al., 2023a) 21.70 31.10 59.77 31.25 41.38 35.96 36.49 -7.23
MoELoRA (Chen et al., 2024a) 17.20 32.70 55.33 32.50 41.41 34.43 35.36 -9.30
MoELoRA (Chen et al., 2024a) 13.90 29.70 54.95 32.50 41.91 32.76 35.66 -12.20
HiDe (Guo et al., 2025a) 17.70 33.00 41.12 20.28 37.27 28.02 33.25 -12.33
HiDe (Guo et al., 2025a) 25.30 29.20 42.13 20.28 35.93 29.23 33.39 -8.93
DISCO (Guo et al., 2025b) 30.60 33.10 65.36 27.50 39.21 39.14 36.73 -0.10
DISCO (Guo et al., 2025b) 32.30 32.30 64.97 30.14 40.46 39.93 37.63 -0.71
MR-LoRA (Ours) 33.00 35.70 67.51 33.75 42.56 42.49 38.85 -0.09
@@ 5.2 Results and Analysis
Domain Continual Learning. As demonstrated in Tab.˜2 (LLaVA-based) and Tab.˜4 (InternVL-based), our proposed MR-LoRA method achieves state-of-the-art performance on the DCL setting, showcasing its exceptional ability to acquire new domain knowledge while preserving previously learned capabilities. The performance of MR-LoRA highlights several key advantages: (1) Approaching Oracle Performance: Our method’s final accuracy on all individual tasks nearly matches the “Oracle” performance. For instance, in Tab.˜2, the final accuracies of MR-LoRA across the five domains are almost identical to the Oracle scores. This indicates that our MLLM-based router can select the most appropriate expert module for each input sample with high precision, allowing the overall performance to approach the theoretical upper bound of a perfect selection mechanism. (2) Superiority over Existing Baselines: In contrast, other baseline methods exhibit significant performance degradation. Parameter-sharing and regularization methods like LoRA-FT and O-LoRA suffer from severe forgetting, as evidenced by their deeply negative BWT scores (e.g., -14.97 for LoRA-FT on LLaVA). This empirically confirms our hypothesis in Sec.˜4.1 regarding the severe task conflict among heterogeneous domains, where shared parameters compromise existing abilities while learning new ones. Although replay-based methods (marked with ) alleviate forgetting by rehearsing old data, their performance remains far inferior to MR-LoRA. Even more advanced baselines like DISCO and SEFE still show a significant gap compared to ours.
Figure 6: Examples demonstrating that the selected expert handles certain questions better than the original expert in DCL and ACL. MLLM-enhanced router selects the most appropriate experts.
Ability Continual Learning. The effectiveness of our proposed method in the more challenging ACL setting is demonstrated in Tabs.˜3 and 5. This setting evaluates the model’s capacity to acquire fundamental new skills and generalize to non-IID test sets. Firstly, we observe that most baselines suffer from severe catastrophic forgetting, revealing a critical weakness in existing CL approaches when faced with real-world, practical non-IID scenarios. In contrast, our MR-LoRA significantly outperforms all baseline methods and successfully improves performance across all four abilities by isolating abilities into dedicated expert modules and leveraging an intelligent MLLM-based router.
Table 6: Ablation study of LoRA rank for each expert LoRA (LLaVA, DCL, last accuracy).
@@
8 80.96 64.64 54.00 55.44 90.75
16 80.92 65.11 53.98 55.90 91.02
32 80.87 65.32 54.12 56.71 91.12
64 81.18 66.07 54.31 56.90 91.60
128 81.14 66.49 54.00 57.63 91.44
Interestingly, the results also reveal a knowledge transfer enabled by our MLLM-enhanced router. In the InternVL experiments, the final accuracy of MR-LoRA on the OCR task is 33.00%, which is higher than the 32.20% achieved by the Oracle. This suggests that the router’s flexible selection mechanism can sometimes leverage knowledge from other related experts (e.g., using the OCR capabilities in the M & L expert) to achieve a result superior to that of a single, isolated specialist. This phenomenon highlights the rationality and sophisticated decision-making capability of the MR-LoRA framework. Fig.˜6 shows the knowledge transfer phenomenon in DCL and ACL.
Rank of Expert LoRA. From the results in Tab.˜6, we find that our method performs well even at very low ranks (e.g., 8), demonstrating its parameter efficiency. This indicates that even if the number of tasks to be learned is large, our method can still achieve a good performance with only a small increase in parameters. Besides, as the expert rank increases, performance can be improved slightly because of more trainable parameters.
Router Accuracy. We ablate the number of samples for routing data and report the router selection accuracy and the last accuracy in domain and ability continual learning. The results are shown in Secs.˜5.2 and 8. In DCL, we find that our method can achieve an excellent performance (almost 100% selection accuracy) using only 20 samples to train the router, which means our method closes the gap of training each task individually. Note that the number of samples we used is much smaller than the number of training samples (60k). Besides, with more sampling data, the router selection accuracy improves and the performance of MR-LoRA slightly increases. In ACL, the performance of MR-LoRA achieves satisfactory performance when the shot of router tuning is 10. It is interesting that the router accuracy of the OCR task is around 50%, but our method can achieve a comparable, or even better performance compared with directly finetuning an OCR LoRA expert (33.60%). This means MR-LoRA uses other experts to solve the OCR task, and these experts perform well on these test samples. It is reasonable that OCR is a basic and fundamental ability that the math and GUI Agent experts are also able to extract equations and web texts from the images.
@@
# Replay
Samples
Router Accuracy (%) Last Accuracy (%)
RS Med AD Sci Fin RS Med AD Sci Fin
100 99.96 99.16 99.98 98.44 99.99 81.04 65.61 54.16 56.77 91.13
50 99.85 98.69 99.94 98.82 100.00 81.00 65.53 54.14 56.76 91.14
30 99.62 98.89 100.00 96.90 99.86 80.92 65.53 54.17 56.59 91.08
20 99.52 97.87 99.89 98.40 99.80 80.87 65.32 54.12 56.71 91.12
10 99.93 98.24 99.93 97.75 99.40 81.04 65.40 54.16 56.63 91.01
Table 8: Router accuracy under different amount of replay data in ability continual learning.
# Replay
Samples
Router Accuracy (%) Last Accuracy (%)
OCR M&L VP GUI Agent OCR M&L VP GUI Agent
100 72.10 94.60 99.87 100.00 32.80 36.30 65.10 32.50
50 65.30 83.90 99.11 100.00 32.70 36.10 64.85 32.50
30 53.60 90.90 97.21 98.38 33.80 36.70 64.85 32.50
20 51.40 86.00 100.00 100.00 33.70 36.20 65.10 32.50
10 81.90 76.30 100.00 100.00 32.80 35.80 65.10 32.50
@@ 6 Conclusion
In this paper, we first propose MLLM-CL benchmark, a novel benchmark including domain continual learning and ability continual learning. In domain continual learning, we select five specific domains (remote sensing, medical, science, autonomous driving, and finance) and focus on IID evaluation. In ability continual learning, we consider a more practical setting where the training and test sets are non-IID. We select four common and fundamental abilities for MLLM to learning sequentially: OCR, math & logic, visual perception, and GUI agent. To solve the two settings in the MLLM-CL benchmark, we first analyze the task conflict between different tasks and then propose an MLLM enhanced router selection method MR-LoRA. Comprehensive experiments and analyses validate the necessity of our MLLM-CL benchmark and show the effectiveness and efficiency of our proposed method. We believe that our carefully designed benchmark and MR-LoRA can serve as a foundation for continual learning in multimodal large language models and will introduce an innovative and practical direction of continual learning and MLLM to the community.
Ethics statement
Our research is grounded in ethical practices, with particular attention paid to the responsible use of data. This work exclusively employs public, well-established datasets from the MLLM community, and we list all used assets’ licenses in Tab.˜12. Our use of this data is in accordance with their provided licenses and intended academic purpose.
Reproducibility statement
To facilitate the reproducibility of our research, we provide comprehensive implementation details in Appendix˜A, including training procedures and hyperparameters. We also report all the result matrices in Appendix˜C. All source code, datasets, and trained models will be publicly released upon the paper’s acceptance.
Appendix